OpenStudy (anonymous):

10/21= 2a/a+b Find the ratio of A to B

OpenStudy (anonymous):

answer must be 5:16??

OpenStudy (kc_kennylau):

Cross multiplication \(\left(\dfrac ab=\dfrac cd\implies ad=bc\right)\)